# Explore then Determine: A GNN-LLM Synergy Framework for Reasoning over Knowledge Graph

### v1

@@ Abstract
The task of reasoning over Knowledge Graphs (KGs) poses a significant challenge for Large Language Models (LLMs) due to the complex structure and large amounts of irrelevant information. Existing LLM reasoning methods overlook the importance of compositional learning on KG to supply with precise knowledge. Besides, the fine-tuning and frequent interaction with LLMs incur substantial time and resource costs. This paper focuses on the Question Answering over Knowledge Graph (KGQA) task and proposes an Explore-then-Determine (EtD) framework that synergizes LLMs with graph neural networks (GNNs) for reasoning over KGs. The Explore stage employs a lightweight GNN to explore promising candidates and relevant fine-grained knowledge to the questions, while the Determine stage utilizes the explored information to construct a knowledge-enhanced multiple-choice prompt, guiding a frozen LLM to determine the final answer. Extensive experiments on three benchmark KGQA datasets demonstrate that EtD achieves state-of-the-art performance and generates faithful reasoning results.
@@ 1 Introduction
Question Answering over Knowledge Graph (KGQA) is a task that aims at answering questions expressed in natural language from entities within a given knowledge graph (KG). KGs, storing a vast amount of facts in the form of triples (e.g., Wikidata [1], YAGO [2], and NELL [3]), are vital for a variety of applications due to their capacity to deliver explicit knowledge. This importance has led to KGQA gaining significant attention from researchers, given its crucial role in numerous intelligent applications such as Apple Siri and Microsoft Cortana [4]. As a complex and difficult task, KGQA usually requires comprehensive question analysis and multi-hop reasoning on the KG to deduce the correct answers [4].
Classical works [5, 6, 7, 8] (e.g., NSM [7] and UniKGQA [8]), usually retrieves a question-specific graph delivering the information related to the question, and then focus on designing different modules to generate answer based on the retrieved graph. Recently, large language models (LLMs) [9, 10, 11, 12], have shown strong power in natural language understanding and processing tasks. But in KGQA task, which predicts answers from KGs, LLMs struggle to excel due to the lack of factual knowledge and the problem of hallucinations [13, 14]. Many efforts have been made to enhance the reasoning ability of LLMs over KGs. However, they all overlook the importance of compositional learning on the graph, which enables the extraction of useful patterns and relationships to supply with precise knowledge related to the question. The text-retrieval methods (e.g.,KAPING [15]), directly retrieve triplets from KGs to construct prompts for LLMs, which often results in redundant information in retrieving [16]. The step-by-step reasoning methods like StructGPT [17], ToG [18], guide LLMs to reason over KGs across multiple steps. Since LLMs lack the inherent capacity to comprehend graph structures, it is often challenging for them to perform effective topological reasoning on graphs [14]. Furthermore, the usage of LLMs, whether through training, fine-tuning, or frequent interaction, entails significant time and resource costs, especially in the KGQA task that requires multi-hop reasoning on a large KG.
In this paper, we propose an Explore-then-Determine framework, named EtD, that synergizes LLM with GNN for effective, efficient and faithful reasoning over KG. The Explore stage is composed of an LLM-empowered GNN module that explores candidates along with relevant knowledge, and adaptively filters irrelevant information in KG. The Determine stage leverages a frozen LLM to determine the final answer with a knowledge-enhanced multiple-choice prompt, which can be seamlessly integrated with any off-the-shelf LLMs during inference, without the burdensome training and fine-tuning. This approach leverages the GNN to perform compositional learning over the KG, identifying the most promising candidates and fine-grained knowledge, while the LLM aligns the question linguistically with the retrieved information to determine the final answer. Two modules are synergized together to benefit each other, as a result prompting the KGQA task. The contributions are summarized as follows:
  • We propose a novel Explore-then-Determine framework, which explores promising candidates and fine-grained knowledge from KG for LLM’s final answer determination, harnessing both the precision of compositional learning and the prowess of language understanding.
@@
  • We design a lightweight GNN model empowered by LLM for precise and efficient KG exploration, and propose a knowledge-enhanced multiple-choice prompt for faithful answer determination.
@@
  • Extensive experiments on KGQA show that our method effectively combines explicit knowledge on KG with implicit knowledge in LLM, outperforming state-of-the-art methods, and achieving efficient and interpretable reasoning over KG.
@@ 2 Related Work
Large Language Models (LLMs). LLMs [9, 10, 11, 12], through pretraining on massive text data, have attained potent language processing ability. They excel in diverse question-answering tasks and demonstrate proficiency in zero-shot learning [15], generating answers based solely on textual prompts. To further improve their performance, some approaches retrieve relevant samples from the training dataset and meticulously design prompts under few-shot learning to stimulate LLMs’ capabilities, such as CoT [19], Self-Consistency [20], and ToT [21]. However, these methods can not compensate for the lack of knowledge in LLMs, and performance gains are limited when applying CoT methods with smaller models ($\sim$100B parameters)[16].
KGQA. Classical embedding-based methods embed entities on KG and design a scoring function to rank them based on their relevance to the question [22, 23]. But these methods are too shallow to capture high-order connectivity and complex semantics. GNN-based works [5, 24, 7, 8] usually follow a retrieval-and-reasoning paradigm, which first retrieves a relatively small question-related subgraph [25] and then use GNNs [26, 27] on the subgraph to find the answer entities. However, they fall short in the understanding of natural language, potentially resulting in semantically incorrect answers. Furthermore, they do not truly unify the subgraph retrieval and reasoning, still separating them into two distinct stages instead of simultaneous processes.
Recently, many works aim to improve the reasoning ability of LLMs over KG to answer complex questions [14]. Semantic parsing methods use LLMs to generate query languages of the question and then execute them on KGs to obtain answers [28, 29, 30, 31]. But they often require fine-tuning of LLMs or the search for a large number of similar examples as prompts. Moreover, the query languages generated often encounter issues of non-executability, yielding no answers [32]. Retrieval-augmented methods retrieve triples from KGs as extra knowledge for LLMs to output the final answers  [33, 15]. These methods tend to neglect the significance of compositional knowledge for precise reasoning. As a result, the retrieved knowledge often contains excess and irrelevant information [34]. Agent-based methods view LLMs as agents and establish an information interaction mechanism between KG and LLMs to iteratively deduce the answer to the question  [17, 18, 35]. They attempt to guide LLMs to reason on graph, but the depth and breadth of this reasoning are limited as LLMs inherently lack the ability to comprehend graph structures. Additionally, frequent interactions with LLMs are inflexible and entail high costs.
@@ 3 Proposed Method
Figure 1: Illustration of the Explore-then-Determine (EtD) framework that synergizes LLMs with GNNs for reasoning over KGs.
@@ 3.1 Overview
Given a natural language question $q$ and a KG $\mathcal{G}=\{(e_{s},r,e_{o})|e_{s},e_{o}\in\mathcal{V},r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities (nodes) and $\mathcal{R}$ is the set of relation types, the task of Question Answering over Knowledge Graph (KGQA) is to find a function $\mathcal{F}(q,G)$ that predicts the answer entities $e_{a}\in\mathcal{V}$ of $q$ over KG $\mathcal{G}$. For each question $q$, the involved topic entity $e_{q}\in\mathcal{V}$ in KG is given.
As a challenging task, KGQA typically requires precise mining of question-related compositional knowledge (such as a list of ordered triplets) in the KG, and demands text comprehension and matching capabilities between questions and entities in KG. However, LLMs is less effective in acquiring the promising candidate entities and relevant knowledge in the KG. Inspired by how humans make decisions when faced with hard tasks by identifying various possible alternatives before choosing the best one, we propose an explore-then-determine (EtD) framework, which adopts exploration module to provide promising candidates and fine-grained knowledge for LLM, and employs determination module to guide LLM to generate final answer.
The proposed framework is shown in Figure 1, which contains two components: (1) semantic-aware graph exploration; (2) knowledge-enhanced answer determination. For the first part, to precisely perceive the question-related compositional knowledge on the graph, we design an LLM-empowered GNN module to adaptively explore the related candidates and relevant knowledge of the given question from KG, i.e., $(\mathcal{C}_{q},\mathcal{K}_{q})=f_{\text{exp}}(q,\mathcal{G})$. For the second part, to effectively leverage the information explored in the first stage, we carefully design a knowledge-enhanced multiple-choice prompt, guiding the LLM to determine final answer based on both the explicit knowledge from KG and the implicit knowledge inside LLM, i.e., $e_{a}=g_{\text{det}}(q,\mathcal{C}_{q},\mathcal{K}_{q})$.
@@
Figure 2: Illustration of semantic-aware graph exploration. We start exploration from topic entity Birdy. In each step, we firstly get the unpruned candidate set $\mathcal{C}_{q}^{\ell}$, calculate the attention weights $\alpha_{q|sr}^{\ell}$ of different edges, prune several irrelevant entities (in white), and update candidate set $\widetilde{\mathcal{C}}_{q}^{\ell}$ (in blue). The representations are propagated from entities in $\widetilde{\mathcal{C}}_{q}^{\ell-1}$ to $\widetilde{\mathcal{C}}_{q}^{\ell}$ through an one-layer GNN.
3.2 Semantic-Aware Graph Exploration
@@
In this section, we aim to extract the promising candidates and relevant compositional knowledge of given question from KG, such that we can provide the LLM with supportive information at once, avoiding the high costs associated with multiple interactions. Due to a large amount of irrelevant information present in the KG, we design an adaptive propagation GNN that, with the semantic representation capabilities provided by the LLM, can automatically prune irrelevant information in KG, enabling precise, efficient and interpretable reasoning.
@@
Semantic-aware pruning. Given the topic entity $e_{q}$ involved in the question $q$, we aim to gradually explore the potential answers starting from $e_{q}$ on the KG. Instead of directly using LLM to explore the candidates, we design a lightweight neural network empowered by LLM, serving as a graph explorer. Specifically, we initialize the candidate set $\mathcal{C}_{q}^{0}\equiv\{e_{q}\}$. The representation of $e_{q}$ is initialized as the question encoding, i.e., $\bm{h}_{e_{q}}^{0}=\bm{h}_{q}=\bm{W}_{L}\cdot\text{LLM}(q)$, where $\text{LLM}(\cdot)$ computes the average embedding in the last layer of Llama2-13B, and $\bm{W}_{L}\in\mathbb{R}^{d\times d_{L}}$ is a learnable weighting matrix mapping the representation to a lower dimension $d$. The representations of other entities are initialized as $\bm{0}$. Assume in the $\ell$-th step ($\ell=1,2,\ldots,L)$, we have explored a set $\mathcal{C}_{q}^{\ell-1}$ of current candidates. The set is then expanded by propagating to the neighbors of entities in $\mathcal{C}_{q}^{\ell-1}$, resulting in an updated set $\mathcal{C}_{q}^{\ell}=\{e_{o}:(e_{s},r,e_{o})\in\mathcal{G},e_{s}\in\mathcal{C}_{q}^{\ell-1}\}$.
@@
where $\sigma$ is the sigmoid function, $\bm{W}^{\ell}$’s in $\mathbb{R}^{1\times d}$ are learnable weight matrices, and $\odot$ is the Hadamard product of vectors. The representations $\bm{h}_{q}$ and $\bm{h}_{r}$ are textual encodings mapped from an LLM (e.g., Llama2-13B) such that the semantic relevance of question with the current edge can be measured. Note that the LLM used as text encoding will not be updated. The representation $\bm{h}_{s}^{\ell-1}\in\mathbb{R}^{d}$ of head entity $e_{s}$ contains the knowledge learned in the $(\ell-1)$-th step.
@@
where $\delta(\cdot)$ is the activation function, $\widetilde{\mathcal{N}}_{e_{o}}^{\ell}$ is the set of preserved neighbor edges of tail entity $e_{o}$, $\bm{W}^{\ell}\in\mathbb{R}^{d\times d}$ is a learnable weight matrix in the $\ell$-th step, and the attention weight $\alpha_{q|sr}^{\ell}$ is computed in Eq.(1). In this way, the compositional information of edges relevant to question $q$ connecting from topic entity $e_{q}$ to $e_{o}$ can be propagated into $\bm{h}_{o}^{\ell}$.
@@
Through adaptive propagation with GNN encoding, the exploration module can achieve effective compositional learning to identify promising candidates. Moreover, with different edge weights, we can extract the relevant triplets as fine-grained knowledge $\mathcal{K}_{q}$, which are organized into relational paths in Section 3.3, effectively illustrating the reasoning process of our model.

### v2

# Dual Reasoning: A GNN-LLM Collaborative Framework for Knowledge Graph Question Answering

@@ Abstract
Large Language Models (LLMs) excel at intuitive, implicit reasoning. Guiding LLMs to construct thought chains can enhance their deliberate reasoning abilities, but also faces challenges such as hallucination. Knowledge Graphs (KGs) can provide explicit structured knowledge for LLMs to alleviate these issues. However, existing KG-enhanced methods often overlook explicit graph learning, making it challenging to efficiently provide precise reasoning chains for LLMs. Following dual-process theory, we propose Dual-Reasoning (DualR), a novel framework that integrates an external system based on Graph Neural Network (GNN) for explicit reasoning on KGs, complementing the implicit reasoning of LLMs through externalized reasoning chains. DualR designs an LLM-empowered GNN module for explicit learning on KGs, efficiently extracting high-quality reasoning chains. These reasoning chains are then refined to a knowledge-enhanced multiple-choice prompt, guiding a frozen LLM to reason thoughtfully for final answer determination. Extensive experiments on three benchmark KGQA datasets demonstrate that DualR achieves state-of-the-art performance while maintaining high efficiency and interpretability. Our code and data are available in https://github.com/leolouis14/DualR.
@@ 1 Introduction
Large language models (LLMs)  [1, 2, 3, 4] have demonstrated impressive capabilities across various natural language processing tasks. Pre-trained on extensive corpora, LLMs excel at implicit and associative reasoning. To further enhance LLMs’ reasoning abilities, many approaches (e.g., Chain-of-Thought (CoT)[5], Tree-of-Thought (ToT)[6]) guide LLMs to generate intermediate steps and form a complete thought chain, aiming for a more deliberate and explicit reasoning [7, 8]. While these approaches can improve performance, they typically increase inference costs, and encounter the challenge of hallucinations when the model lacks relevant knowledge, especially domain-specific and up-to-date knowledge [7, 9].
Integrating external knowledge sources, such as knowledge graphs (KGs), offers a promising solution to these limitations. KGs, storing a vast amount of facts in the form of triples (e.g., Wikidata [10], YAGO [11], and NELL [12]), are vital for a variety of applications due to their capacity to deliver explicit knowledge [13, 14]. A common and essential task for integrating LLMs with KGs is Question Answering over Knowledge Graph (KGQA), which aims at answering natural language question from entities within a given KG. To accurately respond to a given question, a key challenge is that how to enable LLMs to effectively acquire supportive reasoning evidence from a large and complex KG structure. Existing methods frequently overlook the importance of explicit learning within graph structures, which is essential for supplying precise evidence chain for LLM’s reasoning. The text-based retrieval methods (e.g.,KAPING [15]), directly retrieve triplets based on text similarity from KGs for LLMs, which frequently result in redundant or irrelevant information [16]. Another LLM-based retrieval methods like StructGPT [17], ToG [18], guide LLMs to retrieve over KGs across multiple steps. Since LLMs lack the inherent capacity to comprehend graph structures, it is often challenging for them to perform effective topological reasoning on graphs [9]. Furthermore, frequent interaction with LLMs entails significant time and resource costs, especially in the KGQA task that requires multi-hop reasoning on a large KG.
To effectively integrate LLMs with KGs and address challenges above, we draw on cognitive science for inspiration. Bengio et al. [19] has highlighted connections between deep learning and the cognitive process described in Daniel Kahneman’s Thinking fast and Slow [20]. The dual-process theory in the book posits that human thinking arises from two complementary systems: the intuitive and implicit System1, and the deliberate and explicit System 2 [21, 22]. Building on this, we introduce Dual-Reasoning (DualR), a novel approach that incorporates an external, structured “System 2” to explicitly reason on KGs, extracting valuable reasoning chains to complement the intuitive and implicit “System 1”-like reasoning of LLMs. To efficiently implement this externalized “System 2”, we utilize graph neural networks (GNNs) [23, 24, 25], which are well-suited for learning within complex graph structures. Specifically, we first design an LLM-empowered GNN module to explicitly reason on the KG, extracting high-quality reasoning chains relevant to the question. Then, the identified reasoning chains are refined into a knowledge-enhanced multiple-choice prompt, guiding a frozen LLM to reason thoughtfully for final answer determination. With the collaboration of “System 1” and “System 2”, DualR enables efficient and explicit reasoning over KGs, enhancing the reasoning capabilities of LLMs through extracted reasoning chains, and achieving accurate and faithful results in KGQA task. Furthermore, the framework is designed to integrate seamlessly with any off-the-shelf LLMs with just one-step inference, without requiring extensive interactions or fine-tuning, thus allowing for resource-efficient deployment. The contributions are summarized as follows:
  • Following dual-process theory, we introduce a novel framework, Dual-Reasoning (DualR), which integrates an external, structured “System 2” for deliberate, explicit reasoning on KGs, complementing the implicit reasoning of LLMs through externalized reasoning chains.
@@
  • To implement this framework, we design a lightweight GNN model empowered by LLM for precise and efficient reasoning on KG, extracting high-quality reasoning chains. We further propose a knowledge-enhanced multiple-choice prompt to guide the LLM to reason for final answer.
@@
  • Extensive experiments on KGQA show that our method effectively combines GNN’s structured, explicit learning with LLM’s powerful language understanding, outperforming state-of-the-art methods while maintaining high efficiency and interpretability.
@@ 2 Related Work
Dual-Process Theory. Dual-process theory [21, 22, 20] is a psychological account of how human thinking and decision-making arise from two distinct systems. System 1, corresponding to the implicit process, is associative and intuitive, enabling quick comprehension through associations and pre-existing knowledge [7, 8]. System 2, corresponding to the explicit process, is more deliberate and logical. It operates on symbolic structures, conducting explicit reasoning to arrive at conclusions [22, 8]. These systems serve complementary functions and can collaborate for a reasoning problem [22].
LLM reasoning. LLMs [1, 2, 3, 4], have demonstrated impressive capabilities in many tasks such as question answering by leveraging implicit, associative reasoning similar to the intuitive “System 1” process in dual-process theory [22, 20]. To further improve their performance, some approaches instruct LLMs to generate reasoning process in their outputs [5, 26, 6]. While these “System 2”-like deliberate and explicit reasoning methods can achieve better results, they typically entail higher inference costs [7]. Furthermore, LLMs may make mistakes during the reasoning process, due to limitations in their internal knowledge, leading to issues such as hallucinations and inaccuracies [9].
KGQA. Given a natural language question $q$ and a KG $\mathcal{G}=\{(e_{s},r,e_{o})|e_{s},e_{o}\in\mathcal{V},r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities (nodes) and $\mathcal{R}$ is the set of relation types, the task of Question Answering over Knowledge Graph (KGQA) is to find a function $\mathcal{F}(q,G)$ that predicts the answer entities $e_{a}\in\mathcal{V}$ of $q$ over KG $\mathcal{G}$. As a common and practical setting, for each question $q$, the involved topic entities $e_{q}\in\mathcal{V}$ and answer entities $e_{a}\in\mathcal{V}$ are both labeled in KG . KGQA plays a vital role in various intelligent systems, such as Apple Siri and Microsoft Cortana [27]. To solve this task, classical methods [28, 29, 30, 31] usually first retrieve a question-related subgraph and then use different model (e.g., embedding model, graph neural networks[23, 24]) to reason for answers.
LLM for KGQA. Recently, considering the powerful language processing capabilities of LLMs, many works integrate LLMs with KG for KGQA task. These works typically follow the RAG paradigm [32, 33], which utilize various modules to extract knowledge from KG for LLMs generating final answers. Some methods retrieve information from KG based on text similarity [15, 34, 35]. For instance, KAPING [15] employs embedding model to retrieve triplets, but fails to leverage structured knowledge to do reasoning, resulting in the retrieval of excessive and irrelevant information [36]. Another strategy involves guiding the LLM itself to retrieve the knowledge from KG [17, 18, 37, 38]. For example, StructGPT [17] and ToG [18] view LLM as an agent, guiding it to search on the KG iteratively. But the depth and breadth of this search are limited as LLMs inherently lack the ability to comprehend graph structures [9]. Additionally, frequent interactions with LLMs are inflexible and entail high costs. RoG [38] fine-tunes an LLM to generate reasoning paths for information retrieval, but it requires substantial fine-tuning and may generate invalid path due to the hallucinations. A concurrent approach, GNN-RAG [39], utilizes an off-the-shelf GNN model to retrieve the shortest path to high-scoring candidate entities for LLMs. However, it depends on heuristic technique for path retrieval, which may cause low-quality reasoning chains. Moreover, supplying only the path does not fully leverage the GNN model, limiting its effectiveness. Notably, there is an independent category of methods, semantic parsing methods, which use LLMs to generate query languages and then execute them on KGs to obtain answers [40, 41, 42, 43]. These methods typically rely on SPARQL query and are orthogonal to ours.
@@ 3 Proposed Method
Figure 1: Illustration of the Dual-Reasoning (DualR), a GNN-LLM collaborative framework for knowledge graph question answering. It contains two-tier reasoning: (1) semantic-aware knowledge exploration; and (2) knowledge-enhanced answer determination.
@@ 3.1 Overview
Constructing high-quality reasoning chains is crucial for LLMs to generate correct answers. However, in the KGQA task, LLMs face challenges in effectively and efficiently extracting meaningful reasoning chains relevant to a given question from large and complex KGs. Following dual-process theory that describes human thinking as arising from two complementary systems, we propose a novel Dual-Reasoning (DualR) framework, which combines the strengths of GNN-based “System 2” and LLM-based “System 1” to achieve accurate, efficient and faithful reasoning over KGs.
The proposed framework is shown in Figure 1, which contains two-tier reasoning: (1) semantic-aware knowledge exploration; and (2) knowledge-enhanced answer determination. For the first part, to implement the “System 2” efficiently, we design an LLM-empowered GNN module to explicitly reason on the graph, adaptively exploring fine-grained structured knowledge to construct meaningful reasoning chains. For the second part, to effectively leverage the extracted reasoning chains, we carefully design a knowledge-enhanced multiple-choice prompt, guiding the LLM to perform thoughtful reasoning, achieving accurate and reliable answer determination.
@@
Figure 2: Illustration of semantic-aware knowledge exploration. We start exploration from topic entity Birdy. In each step, we firstly get the unpruned candidate set $\mathcal{C}_{q}^{\ell}$, calculate the attention weights $\alpha_{q|sr}^{\ell}$ of different edges, prune several irrelevant entities (in white), and update candidate set $\widetilde{\mathcal{C}}_{q}^{\ell}$ (in blue). The representations are propagated from entities in $\widetilde{\mathcal{C}}_{q}^{\ell-1}$ to $\widetilde{\mathcal{C}}_{q}^{\ell}$ through an one-layer GNN.
3.2 Semantic-Aware Knowledge Exploration
@@
In this section, we aim to construct a deliberate, structured “System 2”-like module on the KG. Through explicit learning and reasoning, it explores structured knowledge on the graph, extracting meaningful reasoning chains that connect topic entity with promising candidates. Due to a large amount of irrelevant information present in the KG, we design an adaptive exploration GNN module that, with the semantic representation capabilities provided by the LLM, can automatically prune irrelevant information, enabling precise, efficient and interpretable reasoning on KG.
@@
Semantic-aware pruning. Given the topic entity $e_{q}$ involved in the question $q$, we aim to gradually explore the potential answer entities starting from $e_{q}$ on the KG. Instead of directly using LLM to explore the candidates, we design a lightweight neural network empowered by LLM. Specifically, we initialize the candidate set $\mathcal{C}_{q}^{0}\equiv\{e_{q}\}$. The representation of $e_{q}$ is initialized as the question encoding, i.e., $\bm{h}_{e_{q}}^{0}=\bm{h}_{q}=\bm{W}_{L}\cdot\text{LLM}(q)$, where $\text{LLM}(\cdot)$ computes the average embedding in the first and last layer of Llama2-13B[3], and $\bm{W}_{L}\in\mathbb{R}^{d\times d_{L}}$ is a learnable weighting matrix mapping the representation to a lower dimension $d$. The representations of other entities are initialized as $\bm{0}$. Assume in the $\ell$-th step ($\ell=1,2,\ldots,L)$, we have explored a set $\mathcal{C}_{q}^{\ell-1}$ of current candidates. The set is then expanded by propagating to the neighbors of entities in $\mathcal{C}_{q}^{\ell-1}$, resulting in an updated set $\mathcal{C}_{q}^{\ell}=\{e_{o}:(e_{s},r,e_{o})\in\mathcal{G},e_{s}\in\mathcal{C}_{q}^{\ell-1}\}$.
@@
where $\sigma$ is the sigmoid function, $\bm{W}^{\ell}$’s in $\mathbb{R}^{1\times d}$ are learnable weight matrices, and $\odot$ is the Hadamard product of vectors. The representations $\bm{h}_{q}$ and $\bm{h}_{r}$ are textual encodings mapped from an LLM (i.e., Llama2-13B) such that the semantic relevance of question with the current edge can be measured. Note that the LLM used as text encoding will not be updated. The representation $\bm{h}_{s}^{\ell-1}\in\mathbb{R}^{d}$ of head entity $e_{s}$ contains the knowledge learned in the $(\ell-1)$-th step.
@@
where $\delta(\cdot)$ is the activation function, $\widetilde{\mathcal{N}}_{e_{o}}^{\ell}$ is the set of preserved neighbor edges of tail entity $e_{o}$, $\bm{W}^{\ell}\in\mathbb{R}^{d\times d}$ is a learnable weight matrix in the $\ell$-th step, and the attention weight $\alpha_{q|sr}^{\ell}$ is computed in (1). In this way, the compositional information of edges relevant to question $q$ connecting from topic entity $e_{q}$ to $e_{o}$ can be propagated into $\bm{h}_{o}^{\ell}$.
@@
Through adaptive propagation with GNN, the “System 2”-like explicit reasoning we build on the KG can effectively explore structured knowledge, identifying promising candidate answers. Furthermore, by leveraging the weights of different edges, we can extract paths connecting the topic entity and candidate answers to construct meaningful reasoning chains, which will be detailed in Section 3.3.